# Partial identification of principal causal effects under violations of principal ignorability

### v1

@@ 7.1 Partial identification related to extended sign issues
A transformation $M$ of random variable $Z$ is said to be an invariant transformation if $M(Z)$ and $Z$ follow the same distribution.
@@
For any invertible $M\in\mathcal{M}(S(1-t)\mid S(t),\boldsymbol{X})$, let $Y^{\prime}(t)=\mu_{yt}(\boldsymbol{X},S(t))+\mu_{yct}(\boldsymbol{X},S(t),M^{-1}(S(1-t)))+\epsilon$. Then, marginally, $(Y^{\prime}(t),S(t))\mid\boldsymbol{X}$ and $(Y(t),S(t))\mid\boldsymbol{X}$ follow the same distribution.
Theorem 3 shows that $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ is unidentifiable, and therefore $E[Y(t)\mid S(t),S(1-t),\boldsymbol{X}]$ and any PCEs of interest are also unidentifiable. Because of the unobserved nature of $S(1-t)$, the flexibility of nonparameteric outcome models allows various possible ways for $S(1-t)$ to contribute to the observed outcome. These plausible outcome models can lead to completely different or even contradictory PCEs. For instance, if $(S(t),S(1-t))$ follows a bivariate Gaussian distribution, then one such invariant transformation can be constructed as $M:s\to 2m_{1-t}(s_{t},\boldsymbol{x})-s$, where $m_{1-t}(s_{t},\boldsymbol{x})$ denotes $E[S(1-t)\mid S(t)=s_{t},\boldsymbol{X}=\boldsymbol{x}]$. Clearly, $M$ is invertible and $M^{-1}=M$. Partial identification arises if $\mu_{yct}(\boldsymbol{x},s_{t},2m_{1-t}(s_{t},\boldsymbol{x})-s_{1-t})\neq\mu_{yct}(\boldsymbol{x},s_{t},s_{1-t})$, which is commonly the case. To provide intuition for this, this inequality would hold in the linear model in (2) if $\beta_{t,1-t}$ is not equal to 0, or equivalently, when principal ignorability fails.
@@ 7.2 Partial identification based on magnitude of unobserved intermediate effect
Under linear outcome models, the sign of the effect of the unobserved intermediate is not identified without additional constraints, but even if the sign were known, the magnitude of the effect is also not identified. Similarly, for nonparametric outcome models, even if the invariant transformations discussed previously were not an issue, we still are not able to identify the magnitude of the effect of the unobserved intermediate variable. In Model (8), this is represented by the magnitude of the $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ function. Conditional on $(S(t),\boldsymbol{X})$,the conditional variance of $\mu_{yct}$ is also the conditional second moment, meaning that its variance can be used to describe the magnitude of $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$. Since $\text{Var}(Y(t)\mid S(t),\boldsymbol{X})=\sigma_{y}^{2}+\text{Var}(\mu_{yct}\mid S(t),\boldsymbol{X})$, and $\text{Var}(Y(t)\mid S(t),\boldsymbol{X})$ is identifiable under standard assumptions, the residual variance $\sigma_{y}^{2}$ can determine the magnitude of $\text{Var}(\mu_{yct}\mid S(t),\boldsymbol{X})$. We can use the residual variance to control the contribution of the unobserved intermediate, with higher $\sigma_{y}^{2}$ indicating less variability of $\mu_{yct}$ and lower $\sigma_{y}^{2}$ indicating greater variability of $\mu_{yct}$. It is not clear what a reasonable bound for $\sigma_{y}^{2}$ is, however, we can use an extension of Assumption 7 to the nonparametric setting to guide this choice. Specifically, we can let $\eta_{A\sim B\mid C}^{2}$ represent nonparametric partial $R^{2}$ values, which are defined as
$\eta_{A\sim B\mid C}^{2}=\dfrac{\text{Var}(E[A\mid B,C])-\text{Var}(E[A\mid C])}{\text{Var}(A)-\text{Var}(E[A\mid C])}.$
@@
$\dfrac{(\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}]))^{2}}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}\leq\sigma_{y}^{2}.$ (9)
@@ 8 Discussion
This paper formalizes inference on principal causal effects when principal ignorability is violated by deriving partial identification results for outcome models when this assumption does not hold, and investigating the implications of this assumption on identification of the crucial association parameter $\rho$. We focused our results on settings with continuous intermediates as this is the scenario that relies most heavily on principal ignorability, though our results also hold for binary intermediates under the same set of assumptions. Our results show that there is an inherent trade-off between the strength of assumptions made and the size of partial identification regions. If principal ignorability, or other assumptions are not made, then all parameters are at best partially identified, and the widths of the partial identification regions for parameters of interest can be exceedingly wide. On the other hand, if strong assumptions are made, then all unknown parameters, including the typically unidentified $\rho$ parameter, can be consistently estimated. We have proposed alternative assumptions that should hold in most applications, which can be applied to sharpen inference and reduce the widths of partial identification regions under mild assumptions. Lastly, many of our results held for linear, parametric outcome models, but we showed that similar ideas can be applied for fully nonparametric outcome models, which are popular in causal inference.
@@
Appendix A Proofs
A.1 Proofs Relating to Model (2)
Consider the following reparameterization for the outcome model in (1):
@@ A.1.1 Proof of Theorem 1
Then, $q_{n}(\rho)=\sum\limits_{i=1}^{n}(t_{i}l_{i}(\rho\mid 1)+(1-t_{i})l_{i}(\rho\mid 0))+\log(f(\rho))$. Let $n_{0}=\sum\limits_{i=1}^{n}(1-t_{i})$ and $n_{1}=\sum\limits_{i=1}^{n}t_{i}$. By the positivity assumption, $n_{0}\to\infty$ and $n_{1}\to\infty$, as $n\to\infty$. Within the control group ($T_{i}=0$) or the treatment group ($T_{i}=1$), $(Y^{\prime}_{i},S^{\prime}_{i})$ are i.i.d.
@@
Since $q_{n}({\rho})=\sum\limits_{i=1}^{n}\left(t_{i}l_{i}(\rho\mid 1)/n_{0}(n_{0}/n)+(1-t_{i})l_{i}(\rho\mid 0)/n_{1}(n_{1}/n)\right)+\log(f(\rho))/n$, the above provides intuition that $\hat{\rho}_{B}$ (posterior mode) converges to ${{\rho}^{\ast}}$.
To complete our proof rigorously, we will show that for any $\epsilon>0$, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$ as $n\to\infty$. If $f(\rho)$ is continuous or bounded, then $\log(f(\rho))/n\to 0$ uniformly as $n\to\infty$. Since $\sum\limits_{t_{i}=t}l_{i}(\rho\mid t)/n_{t}$ uniformly converges to $E(l_{i}(\rho\mid t))$, we have $|q_{n}({\rho})-E(l_{i}(\rho\mid 0))(n_{0}/n)-E(l_{i}(\rho\mid 1))(n_{1}/n)|\to 0$ uniformly in probability, and $P(q_{n}(\rho)\geq q_{n}({{\rho}^{\ast}}):|\rho-{{\rho}^{\ast}}|>\epsilon)\to 0$, that is, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$.
A.1.2 Proof of Theorem 2
@@
We inherit the notations from the proof of Theorem 1. Throughout the proof, we always condition on $\boldsymbol{X}$ and $T$, unless specified. In here we using the Laplace approximation to obtain the asymptotic approximation of posterior variance. Note that the Bernstein–von Mises Theorem provide the justifications for the Laplace approximation.
@@
Therefore, an approximation of posterior variance is $-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}$.
By the Strong Law of Large Number (SLLN), we have
@@
$\begin{split}Var(\rho\mid\mathcal{D})&\sim n^{-1}\left(\bar{T}E\bigg{(}(\frac{\partial l_{i}(\rho|1)}{\partial\rho})^{2}\bigg{)}+(1-\bar{T})E\bigg{(}(\frac{\partial l_{i}(\rho|0)}{\partial\rho})^{2}\bigg{)}\right)^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}\\ &\sim n^{-1}\bigg{(}\bar{T}\beta_{10}^{2}\sigma_{s0}^{2}\bigg{(}\frac{2\rho^{2}\beta_{10}^{2}\sigma_{s0}^{2}}{(\zeta_{1}-\psi_{1}^{2})^{2}}+\frac{1}{\zeta_{1}-\psi_{1}^{2}}\bigg{)}+\\ &\qquad(1-\bar{T})\beta_{01}^{2}\sigma_{s1}^{2}\bigg{(}\frac{2\rho^{2}\beta_{01}^{2}\sigma_{s1}^{2}}{(\zeta_{0}-\psi_{0}^{2})^{2}}+\frac{1}{\zeta_{0}-\psi_{0}^{2}}\bigg{)}\bigg{)}^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}.\end{split}$ (12)
@@ A.1.3 Proof of Proposition 1
In (14), except for signs, there is only one unknown parameter $\sigma_{y}^{2}$.
Since $\beta_{t,1-t}^{2}\geq 0$, (14) also implies that $0\leq\sigma_{y}^{2}\leq\min\limits_{t=0,1}(\zeta_{t}-\psi_{t}^{2})$. Plugging this in first equation of (14), we have
@@
Since $\sigma_{s1}^{2}\beta_{01}^{2}\leq\sigma_{s0}^{2}\beta_{10}^{2}$ (it is equivalent to $\text{Var}(Y(1)\mid S(1),\boldsymbol{X})\geq\text{Var}(Y(0)\mid S(0),\boldsymbol{X})$), then it can be simplified as
@@ A.1.6 Proof of Proposition 3
Combing it with the proof for Proposition 1, the corresponding partial identification region can be obtained.
@@
A.2 Proofs Relating to Model (8)
A.2.1 Proof of Theorem 3
@@
Although $M$ could relate to $S(t)$ and $\boldsymbol{X}$, we do not include them in the $M$ for simplicity of notation. Let $S^{\prime}(1-t)$ denote $M(S(1-t))$ and $p_{\epsilon}$ denote the pdf of $\epsilon$. The the conditional distribution of $Y^{\prime}(t)$ given $S(t)$ and $\boldsymbol{X}$ is
@@
A.2.2 Proof of an equivalent expression for Assumption 8
@@
Appendix B Details of MCMC sampling

### v2

@@ 7.1 Partial identification related to extended sign issues
A transformation $M$ of a random variable $Z$ is said to be an invariant transformation if $M(Z)$ and $Z$ follow the same distribution.
@@
For any invertible $M\in\mathcal{M}(S(1-t)\mid S(t),\boldsymbol{X})$, let $Y^{\prime}(t)=\mu_{yt}\{\boldsymbol{X},S(t)\}+\mu_{yct}[\boldsymbol{X},S(t),M^{-1}\{S(1-t)\}]+\epsilon$. Then, marginally, $(Y^{\prime}(t),S(t))\mid\boldsymbol{X}$ and $(Y(t),S(t))\mid\boldsymbol{X}$ follow the same distribution.
Theorem 3 shows that $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ is unidentifiable, and therefore $E\{Y(t)\mid S(t),S(1-t),\boldsymbol{X}\}$ and any PCEs of interest are also unidentifiable. Because of the unobserved nature of $S(1-t)$, the flexibility of nonparameteric outcome models allows various possible ways for $S(1-t)$ to contribute to the observed outcome. These plausible outcome models can lead to completely different or even contradictory PCEs. For instance, if $(S(t),S(1-t))$ follows a bivariate Gaussian distribution, then one such invariant transformation can be constructed as $M:s\to 2m_{1-t}(s_{t},\boldsymbol{x})-s$, where $m_{1-t}(s_{t},\boldsymbol{x})$ denotes $E\{S(1-t)\mid S(t)=s_{t},\boldsymbol{X}=\boldsymbol{x}\}$. Clearly, $M$ is invertible and $M^{-1}=M$. Partial identification arises if $\mu_{yct}\{\boldsymbol{x},s_{t},2m_{1-t}(s_{t},\boldsymbol{x})-s_{1-t}\}\neq\mu_{yct}(\boldsymbol{x},s_{t},s_{1-t})$, which is typically the case. To provide intuition for this, this inequality would hold in the linear model in (1) if $\beta_{t,1-t}$ is not equal to 0, or equivalently, when principal ignorability fails.
@@ 7.2 Partial identification based on magnitude of unobserved intermediate effect
Under linear outcome models, the sign of the effect of the unobserved intermediate is not identified without additional constraints, but even if the sign were known, the magnitude of the effect is also not identified. Similarly, for nonparametric outcome models, even if the invariant transformations discussed previously were not an issue, we still are not able to identify the magnitude of the effect of the unobserved intermediate variable. In Model (6), this is represented by the magnitude of the $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ function. Conditional on $(S(t),\boldsymbol{X})$,the conditional variance of $\mu_{yct}$ is also the conditional second moment, meaning that its variance can be used to describe the magnitude of $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$. Since $\text{Var}\{Y(t)\mid S(t),\boldsymbol{X}\}=\sigma_{y}^{2}+\text{Var}\{\mu_{yct}\mid S(t),\boldsymbol{X}\}$, and $\text{Var}\{Y(t)\mid S(t),\boldsymbol{X}\}$ is identifiable under standard assumptions, the residual variance $\sigma_{y}^{2}$ can determine the magnitude of $\text{Var}\{\mu_{yct}\mid S(t),\boldsymbol{X}\}$. We can use the residual variance to control the contribution of the unobserved intermediate, with higher $\sigma_{y}^{2}$ indicating less variability of $\mu_{yct}$ and lower $\sigma_{y}^{2}$ indicating greater variability of $\mu_{yct}$. It is not clear what a reasonable bound for $\sigma_{y}^{2}$ is, however, we can use an extension of Assumption 7 to the nonparametric setting to guide this choice. Specifically, we can let $\eta_{A\sim B\mid C}^{2}$ represent nonparametric partial $R^{2}$ values, which are defined as
$\eta_{A\sim B\mid C}^{2}=\dfrac{\text{Var}\{E(A\mid B,C)\}-\text{Var}\{E(A\mid C)\}}{\text{Var}(A)-\text{Var}\{E(A\mid C)\}}.$
@@
$\dfrac{(\text{Var}\{Y(t)\}-\text{Var}[E\{Y(t)\mid S(t),\boldsymbol{X}\}])^{2}}{\text{Var}\{Y(t)\}-\text{Var}[E\{Y(t)\mid\boldsymbol{X}\}]}\leq\sigma_{y}^{2}.$
@@ 8 Discussion
This paper formalizes inference on principal causal effects when principal ignorability is violated by deriving partial identification results for outcome models when this assumption does not hold, and investigating the implications of this assumption on identification of the crucial association parameter $\rho$. We focused our results on settings with continuous intermediates as this is the scenario that relies most heavily on principal ignorability, though we discuss results for binary intermediates in Appendix C. Our results show that there is an inherent trade-off between the strength of assumptions made and the size of partial identification regions. If principal ignorability, or other assumptions, are not made, then all parameters are at best partially identified, and the widths of the partial identification regions for parameters of interest can be exceedingly wide. On the other hand, if strong assumptions are made, then all unknown parameters, including the typically unidentified $\rho$ parameter, can be consistently estimated. We have proposed alternative assumptions that should hold in most applications, which can be applied to sharpen inference and reduce the widths of partial identification regions under mild assumptions. Lastly, many of our results held for linear, parametric outcome models, but we showed that similar ideas can be applied for fully nonparametric outcome models, which are popular in causal inference.
@@
.1 Proofs Relating to Model (1)
Consider the following reparameterization for the outcome model in (3.1):
@@
Then, $q_{n}(\rho)=\sum\limits_{i=1}^{n}(t_{i}l_{i}(\rho\mid 1)+(1-t_{i})l_{i}(\rho\mid 0))+\log(f(\rho))/n$. Let $n_{0}=\sum\limits_{i=1}^{n}(1-t_{i})$ and $n_{1}=\sum\limits_{i=1}^{n}t_{i}$. By the positivity assumption, $n_{0}\to\infty$ and $n_{1}\to\infty$, as $n\to\infty$. Within the control group ($T_{i}=0$) or the treatment group ($T_{i}=1$), $(Y^{\prime}_{i},S^{\prime}_{i})$ are i.i.d.
@@
Since $q_{n}({\rho})=\sum\limits_{i=1}^{n}\left(t_{i}l_{i}(\rho\mid 1)/n_{0}(n_{0}/n)+(1-t_{i})l_{i}(\rho\mid 0)/n_{1}(n_{1}/n)\right)+\log(f(\rho))/n$, the above provides intuition that $\hat{\rho}_{B}$ (posterior mode) converges to ${{\rho}^{\ast}}$. To complete our proof rigorously, we will show that for any $\epsilon>0$, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$ as $n\to\infty$. If $f(\rho)$ is continuous or bounded, then $\log(f(\rho))/n\to 0$ uniformly as $n\to\infty$. Since $\sum\limits_{t_{i}=t}l_{i}(\rho\mid t)/n_{t}$ uniformly converges to $E(l_{i}(\rho\mid t))$, we have $|q_{n}({\rho})-E(l_{i}(\rho\mid 0))(n_{0}/n)-E(l_{i}(\rho\mid 1))(n_{1}/n)|\to 0$ uniformly in probability, and $P(q_{n}(\rho)\geq q_{n}({{\rho}^{\ast}}):|\rho-{{\rho}^{\ast}}|>\epsilon)\to 0$, that is, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$.
.1.2 Proof of Theorem 2
@@
We inherit the notations from the proof of Theorem 1. Throughout the proof, we always condition on $\boldsymbol{X}$ and $T$, unless specified. In here we use the Laplace approximation to obtain the asymptotic approximation of posterior variance. Note that the Bernstein–von Mises Theorem provide the justifications for the Laplace approximation.
@@
Therefore, an approximation of the posterior variance is $-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}$. By the Strong Law of Large Numbers (SLLN), we have
@@
$\begin{split}Var(\rho\mid\mathcal{D})&\sim n^{-1}\left(\bar{T}E\bigg{(}\Big{(}\frac{\partial l_{i}(\rho|1)}{\partial\rho}\Big{)}^{2}\bigg{)}+(1-\bar{T})E\bigg{(}\Big{(}\frac{\partial l_{i}(\rho|0)}{\partial\rho}\Big{)}^{2}\bigg{)}\right)^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}\\ &\sim n^{-1}\bigg{(}\bar{T}\beta_{10}^{2}\sigma_{s0}^{2}\bigg{(}\frac{2\rho^{2}\beta_{10}^{2}\sigma_{s0}^{2}}{(\zeta_{1}-\psi_{1}^{2})^{2}}+\frac{1}{\zeta_{1}-\psi_{1}^{2}}\bigg{)}+\\ &\qquad(1-\bar{T})\beta_{01}^{2}\sigma_{s1}^{2}\bigg{(}\frac{2\rho^{2}\beta_{01}^{2}\sigma_{s1}^{2}}{(\zeta_{0}-\psi_{0}^{2})^{2}}+\frac{1}{\zeta_{0}-\psi_{0}^{2}}\bigg{)}\bigg{)}^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}.\end{split}$ (9)
@@
In (11), except for signs, there is only one unknown parameter $\sigma_{y}^{2}$. Since $\beta_{t,1-t}^{2}\geq 0$, (11) also implies that $0\leq\sigma_{y}^{2}\leq\min\limits_{t=0,1}(\zeta_{t}-\psi_{t}^{2})$. Plugging this into the first equation of (11), we have
@@
Since $\sigma_{s1}^{2}\beta_{01}^{2}\leq\sigma_{s0}^{2}\beta_{10}^{2}$ (which is equivalent to $\text{Var}(Y(1)\mid S(1),\boldsymbol{X})\geq\text{Var}(Y(0)\mid S(0),\boldsymbol{X})$), then it can be simplified as
@@
Combing it with the proof for Proposition 1, the corresponding partial identification region can be obtained. ∎
@@
.2 Proofs Relating to Model (6)
.2.1 Proof of Theorem 3
@@
Although $M$ could relate to $S(t)$ and $\boldsymbol{X}$, we do not include them in the $M$ for simplicity of notation. Let $S^{\prime}(1-t)$ denote $M(S(1-t))$ and $p_{\epsilon}$ denote the pdf of $\epsilon$. Throughout, for simplicity of notation we let $p\big{(}Y^{\prime}(t)=y)$ denote the pdf of $Y^{\prime}(t)$ evaluated at $y$, with similar notation for $S(t)$. The the conditional distribution of $Y^{\prime}(t)$ given $S(t)$ and $\boldsymbol{X}$ is
@@
.2.2 Proof of an equivalent expression for Assumption 8
@@
Appendix A Details of MCMC Sampling